# Synthetic Multimodal Question Generation

### v1

@@ Abstract
Multimodal Retrieval Augmented Generation (MMRAG) is a powerful approach to question-answering over multimodal documents. A key challenge with evaluating MMRAG is the paucity of high-quality datasets matching the question styles and modalities of interest. In light of this, we propose SMMQG, a synthetic data generation framework. SMMQG leverages interplay between a retriever, large language model (LLM) and large multimodal model (LMM) to generate question and answer pairs directly from multimodal documents, with the questions conforming to specified styles and modalities. We use SMMQG to generate an MMRAG dataset of 1024 questions over Wikipedia documents and evaluate state-of-the-art models using it, revealing insights into model performance that are attainable only through style- and modality-specific evaluation data. Next, we measure the quality of data produced by SMMQG via a human study. We find that the quality of our synthetic data is on par with the quality of the crowdsourced benchmark MMQA and that downstream evaluation results using both datasets strongly concur.
@@
C3 AI   Carnegie Mellon University {ian.wu, sravan.jayanthi, simon.rosenberg, sina.pakazad}@c3.ai {sherryw, gneubig}@cs.cmu.edu, vijayv@andrew.cmu.edu
@@ 1 Introduction
To address this issue, we introduce a method for Synthetic Multimodal Question Generation we call SMMQG. SMMQG is a synthetic data generation framework that leverages interplay between a retriever, an LLM and an LMM with in-context learning Brown et al. (2020) to generate multimodal questions and answers based directly on input documents. Crucially, SMMQG enables fine-grained control over the styles and modalities of questions, and is capable of producing both unimodal and cross-modal questions. Using SMMQG, we create a dataset of 1024 questions and answers of various styles and modalities from Wikipedia documents. We then use this dataset to unearth style- and modality-specific insights into the MMRAG performance of various models.
One concern with synthetic data generation is that the resulting data is of low quality. To assuage this concern, we conduct a human study to measure the quality of our dataset. We find that our dataset’s quality is on par with or better than that of popular crowdsourced benchmark MMQA Talmor et al. (2021) when measured along five different metrics. We also show that downstream evaluation results obtained using our SMMQG dataset display strong concurrence Liu et al. (2023b) with those obtained using MMQA, demonstrating that our synthetic dataset can be used in place of MMQA for model selection.
@@ 4.1 Building a Synthetic Multimodal Wikipedia QA Dataset
We use SMMQG to build a QA dataset over Wikipedia documents. We use the text passages, tables and images gathered by Talmor et al. (2021) for MMQA as our sources. These sources include approximately 57,000 captioned images, 232,000 text passages and 12,000 tables, and were scraped from the 2020-01-01 English Wikipedia dump.
We preprocess the dataset by removing text passages with lengths of less than 200 characters, as we find that short passages rarely contain enough information to generate good questions. We generate a total of 1024 QA samples across five different question styles covering all pairwise modality combinations. [^3] SMMQG can be used to generate multimodal questions with three or more question sources Luo et al. (2023). We leave exploration of this to future work. These are summarised in Table 1, and further examples are detailed in Appendix A. We choose these five question styles because (1) they test for a diverse set of reasoning abilities (2) they are well-represented across the QA literature Khashabi et al. (2018); Rajpurkar et al. (2016); Yang et al. (2018); Dua et al. (2019). We plan on releasing our SMMQG-generated dataset following publication of this work.
@@ 4.2 Retriever and QA Model Evaluation
We evaluate the LLM + LMM combinations by conditioning answer generation on the SMMQG questions and question sources [^5] This evaluation setup is easier than end-to-end evaluation where the retrieved sources are used for QA, but has the advantage of allowing for independent evaluation of the retrieval and QA components. and then scoring the predictions against the answers using GPT-4-Turbo as a judge, which has been shown to produce judgements that correlate strongly with human judgements Zheng et al. (2023); Kim et al. (2024). The judge is asked to score the predicted answer on a three-point scale given the SMMQG-generated answer and the question sources. For QA, we use the LMM when the question sources contain at least one image; otherwise we use the LLM. An example of the judge prompt is provided in Appendix F.
We assess the open-source model combinations Vicuna-7b-v1.5 + LLaVA-v1.5-7b, its 13b counterpart Zheng et al. (2023); Liu et al. (2023a), and Qwen-Chat + Qwen-Chat-VL Bai et al. (2023a, b). The LLaVA-v1.5 models are LMMs finetuned from Vicuna-v1.5 Peng et al. (2023), and both Qwen models are 7B parameter models finetuned from Qwen-LM Bai et al. (2023a). We also assess the proprietary multimodal models GPT-4-Turbo OpenAI (2024), Gemini Pro 1.0 Google (2024) and the Claude 3 family Anthropic (2024), which are trained to process text-only inputs in addition to multimodal inputs and so can act as both the LLM and LMM. See Appendix G for further details.
Evaluation Results Retriever evaluation results are shown in the top subtables of Tables 2 and 3. We find that the text-based E5 retriever outperforms BM25 and OpenCLIP across all question styles, and is especially strong on abstractive styles (numerical, compare contrast, multi-hop). We also find that OpenCLIP significantly outperforms text-based retrieval for pure image retrieval questions, but underperforms text-based retrieval for all other modalities.
@@
In summary, we demonstrate that SMMQG can generate question style and modality-specific evaluation datasets. Such datasets reveal important insights into the style- and modality-specific strengths and weaknesses of retrievers and QA models that would otherwise remain hidden.
@@ 5.1 MMQA
In the following sections, we assess the quality of our SMMQG dataset and use MMQA Talmor et al. (2021) as a reference to better understand the significance of our results. MMQA is a crowdsourced benchmark constructed over the same documents as ours. It contains both uni- and cross-modal questions, and provides long-form answers.
@@ 5.2 Human Study
We conduct a human study to directly evaluate the quality of our SMMQG dataset. We randomly draw 300 samples from our dataset, along with 60 single modality and 60 compose samples from the train set of MMQA, ensuring even distribution over modalities. We combine the samples and ask crowdworkers to rate them along five metrics:
@@
We compute the average rating for each metric for each dataset and report the results in Table 4. See Appendix I for further details on our human study methodology and for style-specific results. We employ the Mann-Whitney U test [^7] This non-parametric test compares differences between two independent groups when the dependent variable is either ordinal or continuous but not normally distributed. (for Question Fluency) and Fisher’s exact test [^8] This test is used to determine if there are nonrandom associations between two categorical variables. In our case, the variables are dataset (SMMQG/MMQA) and human judgement for a given metric (Yes/No). (for the remaining metrics) to determine statistical significance. Our findings are as follows:
@@ 5.3 Measuring Concurrence
We randomly draw 150 information extraction and 150 multi-hop samples from our SMMQG dataset, evenly distributed over modalities. We also randomly draw 150 single modality and 150 compose samples from MMQA, again distributed evenly over modalities. We then run evaluation using the methodology described in Section 4.2 (see Appendix H for these results) and calculate concurrence by computing Kendall’s tau on the ranked lists of these results. We report our findings in Table 5.
@@ 7 Conclusion
We introduce SMMQG, a framework for generating synthetic multimodal questions and answers grounded in multimodal documents that adhere to user-specified question styles and modalities. We demonstrate that an SMMQG-generated evaluation dataset is able to reveal novel style- and modality-specific insights into the performance of state-of-the-art retrievers, LLMs and LMMs. Through a human study and by measuring dataset concurrence, we also show that the quality of data generated by SMMQG is on-par with the quality of data from crowdsourced benchmark MMQA and that our dataset can be used in place of MMQA for model selection. We hope that this work will enable automatic, large-scale and tailored evaluation of MMRAG systems, thereby facilitating their adoption in practical applications.
@@ Appendix F GPT-4-Turbo Judge
Appendix G Model Details and Licences
All model evaluation experiments were done using greedy decoding. All inference experiments with open-source models were done using 1 A100 GPU.
@@ Appendix G Model Details and Licences
Appendix H Concurrence Study
H.1 Evaluation Results and Discussion
@@
Tables 6 and 7 contain the SMMQG and MMQA evaluation results used in our concurrence study in Section 5.3. We see that information extraction and multi-hop SMMQG questions are in general easier to answer than their MMQA counterparts. One hypothesis is that this results from MMQA questions being less fluent and less answerable, as shown in Section 5.2. To investigate this, we remove all samples with unanswerable questions and with question fluencies of less than 4 from the datasets used in Section 5.3, and replace them with new samples that we manually validate. We re-run evaluation on this new dataset, but find that MMQA questions remain more challenging than SMMQG questions.
@@
Appendix I Human Study
I.1 Additional Results

### v2

@@ Abstract
Multimodal Retrieval Augmented Generation (MMRAG) is a powerful approach to question-answering over multimodal documents. A key challenge with evaluating MMRAG is the paucity of high-quality datasets matching the question styles and modalities of interest. In light of this, we propose SMMQG, a synthetic data generation framework. SMMQG leverages interplay between a retriever, large language model (LLM) and large multimodal model (LMM) to generate question and answer pairs directly from multimodal documents, with the questions conforming to specified styles and modalities. We use SMMQG to generate an MMRAG dataset of 1024 questions over Wikipedia documents and evaluate state-of-the-art models using it, revealing insights into model performance that are attainable only through style- and modality-specific evaluation data. Next, we measure the quality of data produced by SMMQG via a human study. We find that the quality of SMMQG-generated synthetic data is on par with the quality of the crowdsourced benchmark MMQA and that downstream evaluation results using both datasets strongly concur.
@@
C3 AI   Connectly AI   Carnegie Mellon University {ian.wu, simon.rosenberg, sina.pakazad}@c3.ai, sravan@connectly.ai {sherryw, gneubig}@cs.cmu.edu, vijayv@andrew.cmu.edu
@@ 1 Introduction
To address this issue, we introduce a method for Synthetic Multimodal Question Generation we call SMMQG. SMMQG is a synthetic data generation framework that leverages interplay between a retriever, an LLM and an LMM with in-context learning Brown et al. (2020) to generate multimodal questions and answers based directly on input documents. Crucially, SMMQG enables fine-grained control over the styles and modalities of questions, and is capable of producing both unimodal and cross-modal questions. In order to demonstrate the utility of SMMQG, we first use it to create a dataset of 1024 questions and answers of various styles and modalities from Wikipedia documents. We then use this dataset to unearth novel style- and modality-specific insights into the MMRAG performance of various models. Such insights would be difficult to obtain without style- and modality-specific evaluation data, which SMMQG can produce in an automatic and scalable fashion.
One concern with synthetic data generation is that the resulting data is of low quality. To assuage this concern and verify that SMMQG produces high quality data, we conduct a human study to measure the quality of our synthetic dataset. We find that our dataset’s quality is on par with or better than that of popular crowdsourced benchmark MMQA Talmor et al. (2021) when measured along five different metrics. We also show that downstream evaluation results obtained using our SMMQG dataset display strong concurrence Liu et al. (2023b) with those obtained using MMQA, demonstrating that our synthetic dataset can be used in place of MMQA for model selection.
@@ 4.1 Building a Synthetic Multimodal Wikipedia QA Dataset
We use SMMQG to build a QA dataset over Wikipedia documents. We use this dataset to facilitate experiments (1) demonstrating the utility of SMMQG as a framework for generating style- and modality-specific datasets (Section 4) and (2) verifying that SMMQG produces high quality data (Section 5). We use the text passages, tables and images gathered by Talmor et al. (2021) for MMQA as our sources. These sources include approximately 57,000 captioned images, 232,000 text passages and 12,000 tables, and were scraped from the 2020-01-01 English Wikipedia dump.
We preprocess the dataset by removing text passages with lengths of less than 200 characters, as we find that short passages rarely contain enough information to generate good questions. We generate a total of 1024 QA samples across five different question styles covering all pairwise modality combinations. [^3] SMMQG can be used to generate multimodal questions with three or more question sources Luo et al. (2023). We leave exploration of this to future work. These are summarised in Table 1, and further examples are detailed in Appendix A. We choose these five question styles because (1) they test for a diverse set of reasoning abilities (2) they are well-represented across the QA literature Khashabi et al. (2018); Rajpurkar et al. (2016); Yang et al. (2018); Dua et al. (2019).
We also generate an additional dataset from different source documents to demonstrate that SMMQG is compatible with more domain-specific sources. See Appendix J for details.
@@ 4.2 Retriever and QA Model Evaluation
We evaluate the LLM + LMM combinations by conditioning answer generation on the SMMQG questions and question sources [^5] This evaluation setup is easier than end-to-end evaluation where the retrieved sources are used for QA, but has the advantage of allowing for independent evaluation of the retrieval and QA components. and then scoring the predictions against the answers using GPT-4-Turbo as a judge, which has been shown to produce judgements that correlate strongly with human judgements Zheng et al. (2023); Kim et al. (2024). The judge is asked to score the predicted answer on a three-point scale given the SMMQG-generated answer and the question sources. For QA, we use the LMM when the question sources contain at least one image; otherwise we use the LLM. An example of the judge prompt is provided in Appendix F. In addition to GPT-4-Turbo-judge scores, we also report GPT-3.5-Turbo-judge (Brown et al., 2020), BERTScore (Zhang et al., 2020) and ROUGE (Lin, 2004) scores. We include these results in Appendix G.
We assess the open-source model combinations Vicuna-7b-v1.5 + LLaVA-v1.5-7b, its 13b counterpart Zheng et al. (2023); Liu et al. (2023a), and Qwen-Chat + Qwen-Chat-VL Bai et al. (2023a, b). The LLaVA-v1.5 models are LMMs finetuned from Vicuna-v1.5 Peng et al. (2023), and both Qwen models are 7B parameter models finetuned from Qwen-LM Bai et al. (2023a). We also assess the proprietary multimodal models GPT-4-Turbo OpenAI (2024), Gemini Pro 1.0 Google (2024) and the Claude 3 family Anthropic (2024), which are trained to process text-only inputs in addition to multimodal inputs and so can act as both the LLM and LMM. See Appendix I for further details.
Evaluation Results Retriever evaluation results are shown in the top subtables of Tables 2 and 3. We find that the text-based E5 retriever outperforms BM25 and OpenCLIP across all question styles, and is especially strong on abstractive styles (numerical, compare contrast, multi-hop), although we stress that our evaluation process is likely biased towards E5 as it was used in the question generation process itself. We also find that OpenCLIP significantly outperforms text-based retrieval for pure image retrieval questions, but underperforms text-based retrieval for all other modalities. Our findings underscore the importance of using robust, general-purpose multimodal embedders that are capable of performing consistently across diverse modalities, and we hope that our work can be used to support the development of such models.
@@
In summary, we demonstrate that SMMQG can generate question style- and modality-specific evaluation datasets. Such datasets reveal important insights into the style- and modality-specific strengths and weaknesses of retrievers and QA models that would otherwise remain hidden.
@@ 5.1 MMQA
In the following sections, we assess the quality of our SMMQG-generated dataset and use MMQA Talmor et al. (2021) as a reference to better understand the significance of our results. MMQA is a crowdsourced benchmark constructed over the same documents as ours. It contains both uni- and cross-modal questions, and provides long-form answers.
@@ 5.2 Human Study
We conduct a human study to directly evaluate the quality of SMMQG-produced data. We randomly draw 300 samples from our dataset, along with 60 single modality and 60 compose samples from the train set of MMQA, ensuring even distribution over modalities. We combine the samples and ask crowdworkers to rate them along five metrics:
@@
We compute the average rating for each metric for each dataset and report the results in Table 4. See Appendix L for further details on our human study methodology and for style-specific results. We employ the Mann-Whitney U test [^7] This non-parametric test compares differences between two independent groups when the dependent variable is either ordinal or continuous but not normally distributed. (for Question Fluency) and Fisher’s exact test [^8] This test is used to determine if there are nonrandom associations between two categorical variables. In our case, the variables are dataset (SMMQG/MMQA) and human judgement for a given metric (Yes/No). (for the remaining metrics) to determine statistical significance. Our findings are as follows:
@@ 5.3 Measuring Concurrence
We randomly draw 150 information extraction and 150 multi-hop samples from our SMMQG dataset, evenly distributed over modalities. We also randomly draw 150 single modality and 150 compose samples from MMQA, again distributed evenly over modalities. We then run evaluation using the methodology described in Section 4.2 (see Appendix K for these results) and calculate concurrence by computing Kendall’s tau on the ranked lists of these results. We report our findings in Table 5.
@@ 7 Conclusion
We introduce SMMQG, a framework for generating synthetic multimodal questions and answers grounded in multimodal documents that adhere to user-specified question styles and modalities. We demonstrate that an SMMQG-generated evaluation dataset is able to reveal novel style- and modality-specific insights into the performance of state-of-the-art retrievers, LLMs and LMMs. Through a human study and by measuring dataset concurrence, we also show that the quality of data generated by SMMQG is on-par with the quality of data from crowdsourced benchmark MMQA and that our dataset can be used in place of MMQA for model selection. We hope that SMMQG will enable automatic, large-scale and tailored evaluation of MMRAG systems, thereby facilitating their adoption in practical applications.
@@ Appendix F GPT-4-Turbo Judge
Appendix G Evaluation Results with Additional Metrics
In addition to GPT-4 scores, we also report evaluation results computed using GPT-3.5-Turbo-judge, BERTScore and ROUGE scores. For GPT-3.5-Turbo, we ask the judge to compare the model answer (reference) against the predicted answer given the question and to score the answer as correct, incorrect or partially correct. Our GPT-3.5.-Turbo judge prompt is as follows:
You are a fair judge. You are provided with a question, a reference answer and a prediction. Decide whether the prediction is correct based on the reference answer on a scale of 0 to 2. If the prediction is incorrect, return a 0. If the prediction is partially correct, return a 1. If the prediction is correct, return a 2. Do not use your own knowledge. Use only the reference.
Example 1
Question: Who was the F1 champion in 2019?
Reference: Lewis Hamilton
Prediction: Seb Vettel
Score: 0
Example 2
Question: Explain the purpose of dropout in neural networks.
Reference: Dropout is used for regularisation. It helps prevent overfitting.
Prediction: Reduction in overfitting via regularisation.
Score: 2
Example 3
Question: Compare and contrast the careers of Federer and Nadal.
Reference: Federer has won 20 grand slams, and Nadal has won 22. Federer retired in 2021, whereas Nadal currently still plays tennis.
Prediction: Federer won 20 grand slams and Nadal won 22.
Score: 1
Input
Question: {question}
Reference: {reference}
Prediction: {prediction}
Score:
As GPT-3.5-Turbo lacks multimodal capabilities, we do not provide sources for decision-making - the judge assesses model performance based only on the model answer. For BERTScore, we use microsoft/deberta-xlarge-mnli as the BERTScore model and treat the model answer as the reference and the predicted answer as the candidate. For ROUGE evaluation, we compute ROUGE-1 (unigram overlap), again between the model answer and the predicted answer.
We report aggregate results for each model in Table 6 and report Kendall’s tau on the evaluation ranked lists against GPT-4-Turbo judge in Table 7
GPT-4-Turbo Judge GPT-3.5-Turbo Judge BERTScore ROUGE
Vicuna-7b + LLaVA-7b 45.3 56.9 61.7 53.0
Vicuna-13b + LLaVA-13b 61.3 68.6 69.0 61.3
Qwen-Chat + Qwen-VL-Chat 61.1 69.2 66.8 66.2
Gemini Pro 1.0 73.9 77.1 70.2 63.3
Claude 3 Haiku 62.7 71.5 62.9 71.9
Claude 3 Sonnet 74.1 80.3 66.2 70.9
Claude 3 Opus 86.5 82.7 68.2 72.0
GPT-4-Turbo 93.4 87.8 69.8 77.9
Table 6: Evaluation results with other evaluation metrics. We report results aggregated over all question styles and question modalities.
GPT-3.5-Turbo Judge BERTScore ROUGE
$\tau$ 0.93 0.50 0.72
$p$-value <0.001* 0.11 0.014*
Table 7: Kendall’s tau and its associated $p$-value on the ranked lists of evaluation results against GPT-4-Turbo judge. We find that GPT-3.5-Turbo judge and ROUGE correlate well with GPT-4-Turbo judge.
Appendix H Analysis of Question Style Overlap
In order to better understand how distinct our pre-defined question styles are, we analyse the level of question style overlap in our SMMQG-generated dataset. We use GPT-4-Turbo as a question style binary classifier, pass each QA pair along with descriptions of the question styles to it and ask whether the question belongs to a given style. We define questions with overlap as those that elicit a positive response from the classifier for two or more styles, and report the percentage of questions in our dataset that contain overlap in Table 8, breaking down results by the ground-truth style:
Info Extraction Compare Contrast Numerical Compound Multi-hop
% Overlap 4.3 0.0 4.9 2.8 20.0
Table 8: Percentage of questions with overlap by style. We define questions with overlap as those that are classified by GPT-4-Turbo as belonging to more than one question style. The only question style with significant overlap is the Multi-hop question style.
We see that overlap is insignificant for all but the multi-hop questions. Upon closer inspection, we notice that the multi-hop style commonly overlaps with the numerical and info extraction styles. Our explanation is that multi-hop reasoning inherently requires use of other reasoning skills because these other skills are needed to solve the individual subparts of multi-hop questions.
Appendix I Model Details and Licences
@@
Table 10: Details for open-source LLMs and LMMs.
@@
Table 11: Details for API-based models.
All model evaluation experiments were done using greedy decoding. All inference experiments with open-source models were done using 1 A100 GPU. We document model details and licenses in Tables 9, 10 and 11.
Appendix J Additional Dataset
SMMQG 4.47 94.7 86.4 91.2 91.2
Table 12: Human study results on our additional dataset. These results show that our biology SMMQG dataset quality is high as assessed by crowdworkers.
In addition to our main Wikipedia-derived SMMQG dataset, we also generate a separate QA dataset on a college-level biology textbook (Fowler et al., 2013) containing text and image sources. [^9] https://dept.clcillinois.edu/biodv/PrinciplesOfBiology.pdf We use the exact same SMMQG parameters and prompts as for the original dataset to generate this dataset. Our additional dataset contains 324 questions derived from 862 individual sources.
We conducted an additional human study to assess the quality of this dataset and thereby demonstrate that SMMQG generalises to other, more domain-specific sources than Wikipedia. We sampled 150 questions from our new dataset and asked crowdworkers to assess their quality, repeating the process described in Section 5.2. We report our findings in Table 12. We find that the quality of our new dataset remains high despite the use of a more domain specific dataset. This demonstrates that SMMQG generalises well beyond Wikipedia.
Appendix K Concurrence Study
K.1 Evaluation Results and Discussion
@@
Tables 13 and 14 contain the SMMQG and MMQA evaluation results used in our concurrence study in Section 5.3. We see that information extraction and multi-hop SMMQG questions are in general easier to answer than their MMQA counterparts. One hypothesis is that this results from MMQA questions being less fluent and less answerable, as shown in Section 5.2. To investigate this, we remove all samples with unanswerable questions and with question fluencies of less than 4 from the datasets used in Section 5.3, and replace them with new samples that we manually validate. We re-run evaluation on this new dataset, but find that MMQA questions remain more challenging than SMMQG questions.
@@
Appendix L Human Study
L.1 Additional Results
We present additional human study results breaking down performance by question style in Table 15. We find that SMMQG-question quality is maintained across question styles.